# Rapid and Safe Trajectory Planning over Diverse Scenes through Diffusion Composition

### v1

@@ Abstract
Safe trajectory planning remains a significant challenge in complex environments, where traditional methods often trade off computational efficiency for safety. Comprehensive obstacle modeling improves safety but is computationally expensive, while approximate methods are more efficient but may compromise safety. To address this issue, this paper introduces a rapid and safe trajectory planning framework based on state-based diffusion models. Leveraging only low-dimensional vehicle states, the diffusion models achieve notable inference efficiency while ensuring sufficient collision-free characteristics. By composing diffusion models, the proposed framework can safely generalize across diverse scenarios, planning collision-free trajectories even in unseen scenes. To further ensure the safety of the generated trajectories, an efficient, rule-based safety filter is proposed, which selects optimal trajectories that satisfy both sufficient safety and control feasibility from among candidate trajectories. Both in seen and unseen scenarios, the proposed method achieves efficient inference time while maintaining high safety and stability. Evaluations on the F1TENTH vehicle further demonstrate that the proposed method is practical in real-world applications. The project page is at: https://rstp-comp-diffuser.github.io/.
Index Terms:
Diffusion Model, Composition, Safety, Trajectory Planning, Generalization
{strip}[Uncaptioned image]
Figure 1: Diffusion Composition Enables Efficient, Safe Planning with Practical Real-world Performance. An individual diffusion model cannot ensure safe trajectory planning in out-of-distribution scenarios, whereas composing multiple diffusion models can achieve safety during generalization. Dashed boxes indicate obstacles that do not exist during training. Validation on the F1TENTH platform demonstrates that trajectories planned by the composed diffusion model offer excellent safety while maintaining computational efficiency, indicating that the proposed method is suitable for real-world applications.
@@ I Introduction
Trajectory planning in complex scenes is typically divided into two phases: initial path generation and optimization-based trajectory optimization. Initial collision-free paths can be generated quickly using search or sampling-based methods [1]. However, these methods require significant computational time in complex scenarios [2] and often produce paths with insufficient control feasibility [3]. As a result, optimization-based planning becomes essential to ensure that the planned trajectory is both control-feasible and collision-free, which makes the overall planning process time-consuming. In summary, balancing computational efficiency, control feasibility, and collision avoidance is a challenging yet crucial task for ensuring vehicle safety.
Imitation learning is regarded as an effective solution to this problem, as it can leverage optimization-generated trajectories as expert demonstrations to learn the distribution of optimal trajectories [4]. Among imitation learning methods, the diffusion model is widely used due to its ease of training [5], stable generation process [6], and support for multi-modal inputs [7]. However, the diffusion model suffers from slow trajectory generation due to its iterative denoising process, which can lead to failure in real-world deployment. To address this challenge, we introduce a fast trajectory‐synthesis framework that seamlessly fuses real‐time perception with diffusion‐based generative models. By distilling rich sensor inputs into compact vehicle‐state embeddings, our approach dramatically accelerates the iterative denoising process while fully preserving the diffusion model’s capacity to produce precise, high‐fidelity trajectories.
At the same time, traditional global trajectory planning methods exhibit significant limitations in handling diverse scenarios, requiring substantial adjustments, such as modifications to initial paths and planner parameters. To overcome this limitation, we introduce a rapid, diffusion‐based approach for collision-free trajectory planning that seamlessly adapts to diverse scenarios. It leverages diffusion composition to ensure safe generalization. Specifically, Figures 1 (a) and (b) show diffusion models trained independently on static and dynamic scenes, respectively. By incorporating the dynamic obstacle from scene (b) into the static scene of (a), a new composite scenario is formed. Neither of these models has encountered similar scenarios during training. But individual diffusion models can be composed to generate safe trajectories in this unseen scene, as shown in Figure 1 (c). Our approach enables optimal test-time decision-making to produce safe behaviors, such as accelerating to bypass or decelerating to avoid obstacles. This allows safe trajectory planning for diverse scenes without requiring extra adjustments.
@@
We validate the proposed method on the popular F1TENTH platform [8], which is an efficient testbed for planning methods that integrates perception, planning, and control in a unified hardware system. Experiments on a custom F1TENTH vehicle confirm the superior real-world performance of our method.
@@
In summary, the proposed Rapid and Safe Trajectory Planning (RSTP) framework based on diffusion models has the following contributions:
@@
    Rapid Trajectory Planning with Strong Control Feasibility: The computational load of the denoising process is reduced by using low-dimensional vehicle states from a perception module, reducing the mean computation time of planning to 0.21s in consumer-level hardware. Moreover, a dataset generation method is proposed, using model predictive control (MPC) for trajectory tracking. It creates high-quality trajectory datasets as expert demonstrations for diffusion training.
@@
    Safe Trajectory Planning over Diverse Scenes through Diffusion Composition: Individual models are composed to plan collision-free trajectories in unseen scenarios, addressing the generalization limitations of imitation learning. Additionally, a rule-based safety filter is proposed to select the safest and most feasible trajectory from the candidate trajectory batch generated by diffusion models, which reduces the collision rate by 3.07% with a minor time increment.
@@
    Validation on a Real-World Scaled F1Tenth Vehicle Platform: The proposed method is validated on an F1TENTH vehicle by combining server-based inference with onboard tracking. The simple pure pursuit controller is used to track the planned trajectory, where the F1TENTH vehicle can accurately execute the control inputs without collision, demonstrating strong control feasibility of our planned trajectory and its applicability to real-world scenarios.
This paper is organized as follows: Section II introduces the use of diffusion models to represent the planning problem. Section III provides a detailed description of the dataset generation process, the composition of diffusion models for diverse scenes, and the safety filter. Section IV presents numerical simulations and experiments based on the F1TENTH platform. Finally, Section V concludes the paper by summarizing the key findings and outlining directions for future research.
@@ II-A Path Planning using Potential Field
Collision-free path planning can be achieved using the potential-based approach [10]. In this method, a potential function $\mathcal{U}(\bm{\tau}):\mathbb{R}^{n\times L}\to\mathbb{R}$ is used to assign a potential value to a path $\bm{\tau}$, based on predefined attractive and repulsive potentials. Here, $n$ represents the state dimension and $L$ is the path horizon. Given an initial path $\bm{\tau}$, the collision-free path can be derived by applying gradient descent to the potential function as follows: $\bm{\tau}^{k}_{1:L}=\bm{\tau}^{k-1}_{1:L}-\lambda\nabla_{\bm{\tau}_{1:L}^{k}}\mathcal{U}(\bm{\tau}_{1:L}^{k}),$ where $\bm{\tau}^{k}_{1:L}$ is the path at step $k$, and $\bm{\tau}^{k-1}_{1:L}$ is the path at the previous step, and $\lambda$ is the step size for an iteration. However, using fixed potential functions can lead to corner cases, which makes it challenging to apply this method in more complex scenarios.
@@ II-B Energy-based Models
Energy-based Models are a class of generative models [11]. They can serve as surrogate models for the potential function, represented by $\mathcal{U}(\tau)\sim E_{\theta}(\bm{\tau})$, addressing the limitations of conventional potential-based methods. Given a path variable $\bm{\tau}$, the probability density $p_{\theta}(\bm{\tau})$ of a collision-free path $\bm{\tau}$ is defined as: $p_{\theta}(\bm{\tau})\propto e^{-E_{\theta}(\bm{\tau})}$, where $E_{\theta}(\bm{\tau}):\mathbb{R}^{n\times L}\to\mathbb{R}$ is an energy-based model, typically implemented using a U-Net architecture. Thus, the path planning process can be modified as follows: $\bm{\tau}_{1:L}^{t-1}=\bm{\tau}_{1:L}^{t}-\lambda\nabla_{\bm{\tau}_{1:L}^{t}}E_{\theta}(\bm{\tau}_{1:L}^{t},t,\mathcal{C})+\eta,$ where $\eta\sim\mathcal{N}(\mathbf{0},{\sigma}_{t}^{2}\bm{I})$, with $\lambda$ and $\sigma_{t}$ being specific scaling values. $t$ is the denoising step. And $\mathcal{C}=\{\mathbf{s}_{s},\mathbf{s}_{g},\mathcal{O}\}$ is the set of constraints, which includes the start state and goal state constraints denoted as $\mathbf{s}_{s}$ and $\mathbf{s}_{g}$, as well as the obstacle pose constraints $\mathcal{O}$. In this paper, $\tau=[\mathbf{s}_{1},\cdots,\mathbf{s}_{L}]_{4\times L}$ denotes a discrete planned path. Specifically, $\mathbf{s}=[x,\,y,\,q_{z},\,q_{w}]^{\top}$ with $[x,\,y]^{\top}$ representing the Cartesian position and $[q_{z},\,q_{w}]^{\top}$ the components of the quaternion encoding the yaw angle.
@@ II-C Diffusion Models
$\mathcal{L}_{\mathrm{MSE}}=\|\epsilon-\epsilon_{\theta}(\bm{\tau}^{t}_{1:L},t,\mathcal{C})\|^{2}$
@@
III The Proposed Rapid and Safe Trajectory Planning Method
Our approach proceeds in three stages. First, we describe a comprehensive dataset‐generation procedure. Second, we achieve generalization to unseen scenes by composing the individually trained diffusion models. Finally, a safety filter is applied to guarantee the safety of the final planned trajectory. The entire pipeline is shown in Fig. 3.
III-A Datasets Generation
For collision avoidance in the static scene, global collision-free paths are generated using the iterative collision avoidance (ItCA) method [3]. The ItCA path planning accounts for the smoothness of gear shifting points (GSPs), ensuring the control feasibility of planned paths. Then, an MPC planner is employed for path tracking to produce the final global trajectory. The entire trajectory enriches the static dataset, implicitly encoding velocity through the density of trajectory points within a fixed interval $\Delta t$. Implicit learning of velocity enables the planning of more complex trajectories, exhibiting various behaviors for collision avoidance, such as decelerating to yield or accelerating to bypass.
@@ III-A Datasets Generation
For global collision-free ItCA path planning, starting states $\mathbf{s}_{s}$ are arbitrarily sampled within the feasible region as shown in Figure 3. The goal states $\mathbf{s}_{g}$ are fixed for all candidate points. The velocity of both $\mathbf{s}_{s}$ and $\mathbf{s}_{g}$ are 0. The reference path pose sequence generated by the ItCA method, $\mathbf{p}_{s_{k}}^{\text{ref}}=[x_{s_{k}}^{\text{ref}},\,y_{s_{k}}^{\text{ref}},\,\varphi_{s_{k}}^{\text{ref}}]^{\top}$, is parameterized by the arc length $s_{k}$.
In dynamic scenes, a penalty term based on the distance between the ego vehicle and dynamic obstacles is added to the MPC planner’s cost function to promote collision avoidance: $J_{\textrm{dynamic}}=\alpha\sum_{i=1}^{N_{\mathrm{o}}}({\|\mathbf{p}_{\text{cur}}-\mathbf{o}_{i}\|_{2}})^{-1}$, where $N_{o}$ is the number of dynamic obstacles and $\mathbf{o}_{i}$ is the pose of the $i$-th obstacle. $\alpha$ is the coefficient. $\mathbf{p}_{\text{cur}}$ is the current vehicle pose. It is emphasized that dynamic obstacle avoidance does not consider static obstacles, which may result in collisions with them in the planned trajectory.
@@
[…]
Figure 3: The Overall Framework of the Proposed Rapid and Safe Trajectory Planning (RSTP) Method. In the dataset-generation stage, the ItCA [3] method first creates a static-scene dataset, which then serves as reference paths for producing a dynamic-scene dataset. The outputs of individual diffusion models are composed at each denoising step for unseen scenes.
During the dataset generation process, the MPC in Eq. (1) is used to perform the path tracking while simultaneously recording the trajectory points. To ensure a consistent trajectory horizon $L$ during dataset generation, horizons shorter than $L$ are padded by duplicating the last value, whereas longer trajectories are truncated.
Ultimately, the collision-free demonstration trajectories for the $i$‑th given start and end states is represented as $\bm{\tau}_{i}^{\text{dem}}=[x_{i},y_{i},q_{i,z},q_{i,w}]^{\top}_{4\times L}.$ In dynamic scenarios, the recorded obstacle trajectory is represented as $\mathcal{O}^{\text{single}}_{i}=[x_{i,\text{obst}},y_{i,\text{obst}}]_{2\times L}^{\top}$. The final trajectory dataset is represented as $D=\{{\bm{\tau}_{i}^{\text{dem}},\mathcal{O}^{\text{single}}}\}_{i=1}^{N_{D}},$ where $N_{D}$ is the number of candidate starting points. A key aspect of generating the dataset is ensuring that the time interval $\Delta t$ remains constant, which is especially critical for collision avoidance.
III-B Diffusion Models Composition

### v2

@@ Abstract
Safe trajectory planning in complex environments must balance stringent collision avoidance with real-time efficiency, which is a long-standing challenge in robotics. In this work, we present a diffusion-based trajectory planning framework that is both rapid and safe. First, we introduce a scene-agnostic, MPC-based data generation pipeline that efficiently produces large volumes of kinematically feasible trajectories. Building on this dataset, our integrated diffusion planner maps raw onboard sensor inputs directly to kinematically feasible trajectories, enabling efficient inference while maintaining strong collision avoidance. To generalize to diverse, previously unseen scenarios, we compose diffusion models at test time, enabling safe behavior without additional training. We further propose a lightweight, rule-based safety filter that, from the candidate set, selects the trajectory meeting safety and kinematic-feasibility requirements. Across seen and unseen settings, the proposed method delivers real-time-capable inference with high safety and stability. Experiments on an F1TENTH vehicle demonstrate practicality on real hardware. Project page: https://rstp-comp-diffuser.github.io/.
@@ I Introduction
Traditional trajectory planners for complex environments typically comprise two phases: initial path generation and subsequent trajectory refinement via optimization. While search- and sampling-based methods can produce initial paths [1], their runtime grows rapidly with scene complexity [2], and the resulting paths are often not kinematically feasible [3]. The subsequent optimization step can enforce feasibility and safety, but adds computational burden. For time-critical applications such as planning for autonomous vehicles, the key challenge is to jointly achieve low computational cost, kinematic feasibility, and reliable collision avoidance.
Imitation learning offers an effective alternative to classical two-stage planners by directly learning from optimization-generated trajectories and modeling the distribution of near-optimal solutions [4], thereby avoiding repetitive global search and post hoc optimization. Among imitation learning methods, the diffusion model is widely used due to its ease of training [5], robust generation sampling process [6], and natural support for multi-modal inputs [7]. However, the diffusion model suffers from slow trajectory generation due to its iterative denoising process, which can lead to failure in real-world deployment. In addition, such imitation learning methods heavily rely on the training data – datasets of suboptimal kinematics or limited diversity can easily compromise performance.
To this end, we propose a diffusion-based Rapid and Safe Trajectory Planning (RSTP) framework that includes an efficient MPC-based dataset generation method, an integrated diffusion planning pipeline that maps raw sensor input to executable trajectories, and a real-world compositional novel scene generalization strategy.
The proposed MPC-based dataset generation method uses differential equations to describe system dynamics, thus eliminating extensive reinforcement learning agent training and enabling plug-and-play use across different scenes. As MPC enforces kinematic constraints, the resulting trajectories are kinematically feasible and well-suited as expert demonstrations. Furthermore, to enable responsive diffusion planning in real-time control scenarios, we introduce a fast trajectory‐synthesis framework that seamlessly integrates real‐time perception with diffusion‐based generative models. By fusing low-level vehicle sensory inputs using Simultaneous Localization and Mapping (SLAM) into compact vehicle states, our approach dramatically accelerates the iterative denoising process while fully preserving the diffusion model’s capacity to produce precise, high‐fidelity trajectories.
In addition, traditional global trajectory planning often requires substantial adjustments to handle diverse scenarios. Compared to traditional planning methods, the proposed RSTP framework can seamlessly adapt to diverse scenarios without additional tuning or demonstrations by leveraging test-time diffusion composition. Specifically, Fig. LABEL:fig:vis_abstract (a) and (b) depict two diffusion models trained separately on a static and a dynamic scene. In an unseen composite scenario, for example, where both static obstacles from (a) and a dynamic obstacle from (b) are presented, our method can still generate safe trajectories by composing these two diffusion models. As shown in Fig. LABEL:fig:vis_abstract (c), this test-time diffusion composition yields adaptive behaviors, such as accelerating to bypass obstacles or decelerating to avoid obstacles with no extra model training.
@@
To harden deployment, we propose a lightweight safety filter to further enhance planning safety in real-world applications. We validate the proposed pipeline on the popular F1TENTH platform [8], which is an efficient testbed for planning methods that integrates perception, planning, and control in a unified hardware system. Experiments on a custom F1TENTH vehicle confirm the superior real-world performance of our framework.
@@
In summary, the proposed RSTP framework based on the diffusion model has the following contributions:
@@
    Feasibility-Aware, Data-Driven Trajectory Planning: We introduce an end-to-end training pipeline that includes an energy-based diffusion trajectory planner and a data generation method based on MPC. We demonstrate that such a framework is able to learn the distribution of collision-free, kinematically feasible trajectories, yielding diffusion samples that are safe and directly trackable without post hoc optimization.
@@
    Rapid and Safe Diffusion Planning from Raw Sensor Inputs: By operating in a low-dimensional vehicle-state extracted from perception, RSTP achieves 0.21s mean planning time on commodity hardware. Furthermore, a rule-based safety filter selects an optimal trajectory from a candidate batch, ensuring safety.
@@
    Real-time Diffusion Composition for Flexible Planning in Diverse Scenes: We formulate a composition mechanism that enables generalization to planning in unseen scenarios and demonstrate its effectiveness in time-critical decision-making scenarios with noisy real-world sensory inputs.
  4. 4.
    Validation on a Real-World Scaled F1TENTH Vehicle Platform: The proposed method is validated on an F1TENTH vehicle using server-based inference with simple onboard pure pursuit tracking. The vehicle follows the planned trajectory accurately and without collisions, demonstrating the kinematic feasibility and real-world applicability of our approach.
This paper is organized as follows: Section II introduces the use of diffusion models to represent the planning problem. Section III explains the method for dataset generation. Section IV provides a detailed description of the composition of diffusion models for diverse scenes and the safety filter. Section V presents numerical simulations and experiments based on the F1TENTH platform. Section VI concludes the paper.
@@ II-A Path Planning using Potential Field
Collision-free path planning can be achieved using the potential-based approach [10]. In this method, a potential function $\mathcal{U}(\bm{\tau}):\mathbb{R}^{n\times L}\to\mathbb{R}$ is used to assign a potential value to a path $\bm{\tau}$, based on predefined attractive and repulsive potentials. Here, $n$ represents the state dimension and $L$ is the path horizon. Given an initial path $\bm{\tau}$, the collision-free path can be derived by applying gradient descent to the potential function as follows: $\bm{\tau}^{k}_{1:L}=\bm{\tau}^{k-1}_{1:L}-\lambda\nabla_{\bm{\tau}_{1:L}^{k}}\mathcal{U}(\bm{\tau}_{1:L}^{k}),$ where $\bm{\tau}^{k}_{1:L}$ is the path at step $k$, and $\bm{\tau}^{k-1}_{1:L}$ is the path at the previous step, and $\lambda$ is the step size for an iteration. However, this method leads to corner cases, making it unsuitable for complex scenarios.
@@ II-B Energy-based Models
Energy-based Models are a class of generative models [11]. They can serve as surrogate models for the potential function, represented by $\mathcal{U}(\tau)\sim E_{\theta}(\bm{\tau})$, addressing the limitations of conventional potential-based methods. Given a path variable $\bm{\tau}$, the probability density $p_{\theta}(\bm{\tau})$ of a collision-free path $\bm{\tau}$ is defined as: $p_{\theta}(\bm{\tau})\propto e^{-E_{\theta}(\bm{\tau})}$, where $E_{\theta}(\bm{\tau}):\mathbb{R}^{n\times L}\to\mathbb{R}$ is an energy-based model, typically implemented using a U-Net architecture. Thus, the path planning process can be modified as follows: $\bm{\tau}_{1:L}^{t-1}=\bm{\tau}_{1:L}^{t}-\lambda\nabla_{\bm{\tau}_{1:L}^{t}}E_{\theta}(\bm{\tau}_{1:L}^{t},t,\mathcal{C})+\eta,$ where $\eta\sim\mathcal{N}(\mathbf{0},{\sigma}_{t}^{2}\bm{I})$, with $\lambda$ and $\sigma_{t}$ being specific scaling values. $t$ is the denoising step. And $\mathcal{C}=\{\mathbf{s}_{s},\mathbf{s}_{g},\mathcal{O}\}$ is the set of constraints, which includes the start state and goal state constraints denoted as $\mathbf{s}_{s}$ and $\mathbf{s}_{g}$, as well as the obstacle pose constraints $\mathcal{O}$. In this paper, $\tau=[\mathbf{s}_{1},\cdots,\mathbf{s}_{L}]_{4\times L}$ denotes a discrete planned path. Specifically, $\mathbf{s}=[x,\,y,\,q_{z},\,q_{w}]^{\top}$ with $x,\,y]^{\top}$ representing the Cartesian position and $q_{z},\,q_{w}]^{\top}$ the components of the quaternion encoding the yaw angle.
@@ II-C Diffusion Models
$\mathcal{L}_{\mathrm{MSE}}=\mathbb{E}_{p(\bm{\tau}),\,\epsilon}\|\epsilon-\epsilon_{\theta}(\bm{\tau}^{t}_{1:L},t,\mathcal{C})\|^{2}$
@@
III MPC-based Data Collection
Fig. 3 overviews the RSTP pipeline, which proceeds in three modules. The first module, a comprehensive data collection procedure, is described in detail in this section.
For collision avoidance in the static scene, global collision-free paths are generated using the iterative collision avoidance (ItCA) method [3]. The ItCA path planning accounts for the smoothness of gear shifting points (GSPs), ensuring the kinematic feasibility of planned paths. Then, an MPC planner is employed for path tracking to produce the final global trajectory. The entire trajectory enriches the static dataset, implicitly encoding velocity through the density of trajectory points within a fixed interval $\Delta t$. Implicit learning of velocity enables more complex trajectories, exhibiting behaviors such as decelerating to yield or accelerating to bypass.
@@ III MPC-based Data Collection
For global collision-free ItCA path planning, starting states $\mathbf{s}_{s}$ are arbitrarily sampled within the feasible region. The goal states $\mathbf{s}_{g}$ are fixed for all candidate points. The velocity of both $\mathbf{s}_{s}$ and $\mathbf{s}_{g}$ is 0. The reference path pose sequence generated by the ItCA method, $\mathbf{p}_{s_{k}}^{\text{ref}}=[x_{s_{k}}^{\text{ref}},\,y_{s_{k}}^{\text{ref}},\,\varphi_{s_{k}}^{\text{ref}}]^{\top}$, is parameterized by the arc length $s_{k}$.
In dynamic scenes, a penalty term based on the distance between the ego vehicle and dynamic obstacles is added to the cost function: $J_{\textrm{dynamic}}=\alpha\sum_{i=1}^{N_{\mathrm{o}}}({\|\mathbf{p}_{\text{cur}}-\mathbf{o}_{i}\|_{2}})^{-1}$, where $N_{o}$ is the number of dynamic obstacles and $\mathbf{o}_{i}$ is the pose of the $i$-th obstacle. $\alpha$ is the coefficient. $\mathbf{p}_{\text{cur}}$ is the current vehicle pose. It is emphasized that dynamic obstacle avoidance does not consider static obstacles, which may result in collisions with them in the planned trajectory.
@@
where $\mathbf{A}=[1,\,1,\,1,\,0]^{\top}$, and $N_{p}$ denotes the prediction horizon. $\bm{\zeta}_{\textrm{cur}}$ is the current vehicle state. In dynamic scenes, $\gamma=1$; otherwise, $\gamma=0$. Following the concept of projected velocity in [12, 13], the value of $s_{k}$ is determined by a fixed projection velocity $v_{\text{ref}}$, such that $s_{k}=s_{k-1}+\Delta t\cdot v_{\text{ref}}$, with $s_{0}=s_{\text{cur}}$. $\bm{\zeta}_{\text{max}}$ and $\bm{\zeta}_{\text{min}}$ denote the upper and lower bounds of the state constraints, while $\mathbf{u}_{\text{max}}$ and $\mathbf{u}_{\text{min}}$ are the upper and lower limits of the control inputs, respectively. $T_{s}$ represents the prediction interval. $Q_{1}$, $Q_{2}$, $R_{1}$, and $R_{2}$ are coefficient matrices.
To ensure a consistent trajectory horizon $L$ during dataset generation, horizons shorter than $L$ are padded by duplicating the last value, whereas longer trajectories are truncated.
Ultimately, the collision-free demonstration trajectories for the ii‑th given start and end states is represented as $\bm{\tau}_{i}^{\text{dem}}=[x_{i},y_{i},q_{i,z},q_{i,w}]^{\top}_{4\times L}.$ In dynamic scenarios, the recorded obstacle trajectory is represented as $\mathcal{O}^{\text{single}}_{i}=[x_{i,\text{obst}},y_{i,\text{obst}}]_{2\times L}^{\top}$. The final trajectory dataset is represented as $D=\{{\bm{\tau}_{i}^{\text{dem}},\mathcal{O}^{\text{single}}}\}_{i=1}^{N_{D}},$ where $N_{D}$ is the number of candidate starting points. A key aspect of generating the dataset is ensuring that the time interval $\Delta t$ remains constant, which is especially critical for collision avoidance.
IV The Proposed Rapid and Safe Trajectory Planning Method
This subsection provides a detailed description of the second and third modules in Fig. 3. The second module achieves generalization to unseen scenes by composing the individually trained diffusion models. The third module applies a safety filter to guarantee the safety of the final planned trajectory.
Figure 3: The Overall Framework of the Proposed Rapid and Safe Trajectory Planning (RSTP) Method. Offline Dataset Generation (Left): the ItCA [3] and MPC-based methods provide kinematically feasible trajectory datasets for training. Diffusion Composition (Middle): individual diffusion models can be flexibly composed to tackle novel scenarios not covered in the training data. Online Inference and Control (Right): The ego vehicle performs real-time inference from its current pose, and the safety filter selects the optimal trajectory per cycle for tracking.
IV-A Diffusion Models Composition